# A Longitudinal Randomized Control Study of Companion Chatbot Use: Anthropomorphism and Its Mediating Role on Social Impacts

### v2

@@ Results
Eighty-five percent of participants ($N=155$) completed the full study. Our study had comparable attrition rates (15% attrition, $N$ dropped =28/183=28/183) to other longitudinal research using online samples that reported attrition rates (Kothe & Ling, 2019). Of the dropped participants, 43% had been assigned to the chatbot group, and 57% had been assigned to the control group. Across conditions, 43% of all drops occurred before the day 7 survey, 46% occurred before the day 14 survey (participants completed the day 7 survey but did not complete the day 14 survey), and 11% occurred before the day 21 survey and interview (participants completed the day 14 survey but not the day 21 survey and interview).
@@
Additional Data
As preregistered, analyses were conducted on the participants who completed the full study (85% of participants, $N=155$). Of participants who completed the full study, 86% ($N=133$) also completed the follow-up survey that was sent four weeks after the final survey and interview (completion by condition: chatbot = 65/78; control = 65/77).
@@
Figure 3: Mediation model for desire to socially connect, magnitude of social impact, and anthropomorphism. Paths between variables designated with letters $a$, $b$, and $c$. Path estimates and significance are indicated (** = $p<0.001$, ns = non-significant).
@@
We used non-parametric Wilcoxon rank-sum tests to evaluate differences by condition over time for desire to socially connect, loneliness, social health, magnitude of social impact, and directional social impact scores. Significant differences were found between the chatbot and control groups in their desire to socially connect. Participants in the chatbot condition, versus the control condition, reported higher desire to socially connect when tested on day 7 (chatbot $M=1.95$, control $M=1.46$, $p<0.01$) and 14 (chatbot $M=1.94$, control $M=1.49$, $p<0.05$). However, no difference was found on day 1 or day 21. No significant results emerged for magnitude of social impact, directional social impact, loneliness, or social health. Compared to those who played word games, those who engaged with the companion chatbot did not have significant impacts to their social health, loneliness, or social interactions or relationships.
@@
To test the hypothesis that people with a higher desire to socially connect on day 1 would anthropomorphize the chatbot more (H2), we ran non-parametric Spearman’s correlations with Holm correction for multiple comparisons between participants’ desire to socially connect scores on day 1 and their anthropomorphism of the chatbot over time. These relationships are shown in Figure 1.
Figure 1A shows the day-1 score for the desire to socially connect versus the anthropomorphism scores averaged over time (an average of anthropomorphism scores at day 7, 14, and 21). Desire to socially connect significantly predicted anthropomorphism ($\rho=0.40$, $p<0.01$). People with a higher desire to socially connect on day 1 were more likely to anthropomorphize the chatbot across 21 days of use. In comparison, for the control group, the relationship between desire to socially connect on day 1 and anthropomorphism of the word games was non-significant ($\rho=0.07$, $p=0.57$).
Figure 1B shows the day-1 scores for the desire to socially connect versus the anthropomorphism scores separated by time point (days 7, 14, and 21). The desire to socially connect on day 1 significantly predicted anthropomorphism at all three time points for the chatbot group (day 7: $\rho=0.34$, $p<0.01$, day 14: $\rho=0.45$, $p<0.001$, day 21: $\rho=0.38$, $p<0.01$). These relationships were non-significant for the control group (day 7: $\rho=0.07$, $p=1.0$, day 14: $\rho=0.04$, $p=1.0$, day 21: $\rho=0.07$, $p=1.0$). H2 was supported: people with a higher desire to socially connect on day 1 were more likely to anthropomorphize the chatbot on day 7, 14, and 21. This finding was specific to the chatbot group; people in the control group with a higher desire to socially connect were not more likely to anthropomorphize the word games.
@@
To support H3, people who anthropomorphized the chatbot more should report that the interaction with the chatbot had a greater social impact on them. To test this hypothesis, we ran non-parametric Spearman’s correlations with Holm correction for multiple comparisons between people’s anthropomorphism of the chatbot and the magnitude of social impact over time. The magnitude of social impact score captures the strength of perceived impact regardless of whether the impact was positive or negative (see Methods).
Figure 2A shows the anthropomorphism scores (averaged over days 7, 14, and 21) versus the social impact scores (averaged over days 7, 14, and 21) for the chatbot group. Anthropomorphism significantly predicted social impact ($\rho=0.52$, $p<0.0001$). People who anthropomorphized the chatbot more indicated that the human-chatbot interaction had a greater impact on their social interactions and relationships with family and friends. For the control group, the relationship between anthropomorphism and social impact of the word games was also significant, with a weaker correlation ($\rho=0.33$, $p<0.01$). Those who tended to anthropomorphize the word games also tended to report that the games had a greater social impact on them.
@@
Figure 2B shows three regression lines for the chatbot group. The first shows the anthropomorphism score as measured on day 7 versus the social impact score as measured on day 7; the second shows anthropomorphism versus social impact, both measured on day 14; and the third shows anthropomorphism versus social impact, both measured on day 21. We found that for the chatbot group, for all three regression lines, anthropomorphism significantly predicted social impact (day 7: $\rho=0.43$, $p<0.0001$, day 14: $\rho=0.48$, $p<0.0001$, day 21: $\rho=0.52$, $p<0.0001$). For comparison, for the control group, the only significant relationship between anthropomorphism and social impact occurred on day 7 ($\rho=0.35$, $p<0.01$) and was not significant for day 14 ($\rho=0.20$, $p=0.09$) or day 21 ($\rho=0.23$, $p=0.09$). H3 was supported: higher anthropomorphism of the chatbot correlated with a higher magnitude of social impact of the chatbot.
@@
The foregoing analysis considered only the magnitude of social impacts, not the direction. Did higher anthropomorphism of the chatbot correlate specifically with more positive perceived social impacts, more negative, or did different people tend to swing different directions? We performed an additional analysis using the directional social impact measure. (The individual, raw scores for the two social impact questions were averaged, resulting in a scale that ranged from positive to negative values.) We found that higher anthropomorphism predicted more positive social impacts at all time points for the chatbot group (day 7: $\rho=0.61$, $p<0.0001$, day 14: $\rho=0.67$, $p<0.0001$, day 21: $\rho=0.71$, $p<0.0001$). For comparison, for the control group, the only significant relationship between anthropomorphism and social impact occurred on day 7 and was positive ($\rho=0.35$, $p<0.01$) and was not significant for day 14 ($\rho=0.20$, $p=0.18$) or day 21 ($\rho=0.16$, $p=0.18$). People who anthropomorphized the chatbot more were more likely to report positive impacts to their social interactions and relationships with family and friends.
@@
In addition to asking how participants perceived the impact of the chatbot on their human relationships, we also tested social health through a separate scale. To support H4, people who anthropomorphized the chatbot more should show greater changes to their social health scores. To test this hypothesis, we ran Spearman’s correlations with Holm correction for multiple comparisons. We tested the relationship between the mean anthropomorphism scores for each participant (averaged over days 7, 14, and 21) and the mean social health scores (averaged over days 7, 14, and 21). We also tested the relationship between the anthropomorphism score and the social health score for each of the three time points individually. We found no significant relationships for the chatbot group ($p>0.05$). We also did not find significant relationships for the corresponding analyses in the control group. H4 was not supported.
@@
To support H5, the degree of anthropomorphism of the chatbot should mediate the relationship between people’s desire to socially connect and the magnitude of social impact of interacting with the chatbot. For this analysis, as preregistered, we did not evaluate social health impacts, as those relationships were non-significant in prior analyses. We ran a mediation model with bootstrapped standard errors (1000 resamples) using the lavaan package (Rosseel 2012). Figure 3 depicts the output of the mediation model, including the paths estimates and significance values.
For the chatbot group, results supported a mediation relationship. We found a significant positive association between the desire to socially connect and anthropomorphism ($a=0.54$, $p<0.001$) and between anthropomorphism and social impacts ($b=0.30$, $p<0.01$). The indirect effect of desire to socially connect, via anthropomorphism, on social impacts, was significant ($ab=0.16$, $p<0.05$, 95% CI: [0.03, 0.33]). In contrast, the direct effect of desire to socially connect on social impacts was not significant ($c=0.13$, $p=0.27$). The total effect was significant ($total=0.29$, $p<0.01$, 95% CI: [0.05, 0.48]), providing evidence for partial mediation. Approximately 57% of the total effect was explained by the indirect effect of anthropomorphism. The remaining 43% of the total effect was not explained by anthropomorphism and may be due to direct effects or other unmeasured mediators. These findings suggest that anthropomorphism plays a significant role in linking people’s desire to socially connect to their reported social impacts of interacting with the chatbot. For people with a desire to socially connect to see social impacts from interacting with an AI chatbot, they must perceive it as humanlike.
In comparison, the results for the control condition indicated that the association between the desire to socially connect and anthropomorphism was not significant ($a=0.08$, $p=0.58$). The association between anthropomorphism and magnitude of social impacts was significant ($b=0.31$, $p<0.01$). The indirect effect was not significant ($ab=0.02$, $p=0.63$), nor was the direct effect of desire to socially connect on magnitude of social impacts ($c=0.16$, $p=0.05$). The total effect was also non-significant ($total=0.19$, $p=0.09$). These findings suggest that the mediating role of anthropomorphism is specific to the chatbot condition. H5 was supported.
@@
The best-fit model outputs from these two models are as follows. The stepwise linear regression model (AIC) included 20 predictors and explained 75% of the variance in anthropomorphism of the chatbot (adj $R^{2}=0.75$, $AIC=146.1$, $p<0.0001$). Significant predictors within the model with positive estimates were the desire to socially connect ($b=0.42$, $p<0.001$), beliefs about AI agency ($b=0.21$, $p<0.01$), the General Attitudes Toward Robots Scale (GAToRS) personal-positive scale (P+) ($b=0.23$, $p<0.05$), social health ($b=1.10$, $p<0.001$), age ($b=0.04$, $p<0.01$), mental health vulnerability ($b=0.34$, $p<0.001$), and identifying as having multiple sexual orientations ($b=0.77$, $p<0.05$). Significant predictors within the model with negative estimates were beliefs about AI experience (b=0.36b=-0.36, $p<0.01$), GAToRS society-negative scale (S-) (b=0.25b=-0.25, $p<0.05$), extraversion (b=0.22b=-0.22, $p<0.01$), self-esteem (b=0.38b=-0.38, $p<0.01$), social competence (b=0.33b=-0.33, $p<0.05$), and identifying as lesbian (b=1.13b=-1.13, $p<0.05$) or the sexual orientation of “prefer to self-describe” (b=2.27b=-2.27, $p<0.01$).
@@
Sexual orientation, age, and social variables were not listed within or found to be significant in the LASSO model, indicating that those factors retained limited additional value for generalization.
@@
We then compared the magnitude of social impact scores across time and at each time point between the vulnerable and non-vulnerable subgroups using Wilcoxon rank-sum tests. We found no significant differences across time for the magnitude of social impact (averaged scores for day 7, 14, and 21) between vulnerable and non-vulnerable subgroups. More vulnerable participants did not report greater social impacts than the non-vulnerable group ($p=0.40$). We then evaluated differences at each time point using Holm’s correction for multiple comparisons and did not find significant differences in magnitude of social impact scores between subgroups. Participants in the vulnerable group did not report greater social impacts than those in the non-vulnerable group on day 7, 14, or 21 ($p$ adj for each comparison >0.05>0.05). For comparison, in the control group, participants in the vulnerable group also did not report greater social impacts than those in the non-vulnerable group ($p$ adj for each comparison >0.05>0.05). H6 was not supported.
Attachment and Addiction to Companion Chatbots
We predicted that people assigned to interact with the companion chatbot would not show signs of attachment and addiction to the chatbot (H7). This hypothesis corresponds to a common concern about the psychological impact of chatbots, with evidence both for and against attachment and addiction. Here, we compared attachment and addiction metrics between the chatbot and control conditions. We operationalized attachment and addiction by two metrics measured in the follow-up survey. First, we used the reported engagement in the daily task after the last day of the study (day 21), by means of a binary variable (yes or no). Second, we used the reported duration of time spent doing so. Duration of engagement was open-ended. All open-ended responses were binned into four categories: a few days, 1-2 weeks, 3-4 weeks, and ongoing.
The number of participants who completed the follow-up survey was not statistically different between the two conditions ($\chi^{2}=0$, $p=1$; chatbot: 65 completed, 13 did not; control: 65 completed, 12 did not).
@@
We then evaluated the duration of post-study engagement in the daily task by condition. Of the 14 people who indicated they engaged with the chatbot after the study ended, six (43%) indicated that their engagement was ongoing, none (0%) said they used it for 3-4 weeks, three (21%) said they used it for 1-2 weeks, and five (36%) said they used it for a few days.
Of the 40 people who indicated they engaged with the word games after the study ended, 26 (65%) indicated that their engagement was ongoing, three (8%) said they used it for 3-4 weeks, six (15%) said they used it for 1-2 weeks, and five (13%) said they used it for a few days.
A higher percentage of people in the control condition reported the longest post-task duration of engagement (ongoing engagement) (65%) as compared to the chatbot condition (43%). However, these ratios were not significantly different from one another from a chi-squared approximation with continuity correction ($\chi^{2}=1.3$, $p=0.26$).
H7 was partially supported. People who played the word games were more likely to have engaged in their daily task after the study had ended, as compared to people who interacted with the chatbot. However, people who reported ongoing engagement did not differ between those who played the word games and those who interacted with the chatbot.

### v3

@@ Results
Eighty-five percent of participants ($N=155$) completed the full study. Our study had comparable attrition rates (15% attrition, $N$ dropped =28/183=28/183) to other longitudinal research using online samples that reported attrition rates (Kothe and Ling 2019). Of the dropped participants, 43% had been assigned to the chatbot group, and 57% had been assigned to the control group. Across conditions, 43% of all drops occurred before the day 7 survey, 46% occurred before the day 14 survey (participants completed the day 7 survey but did not complete the day 14 survey), and 11% occurred before the day 21 survey and interview (participants completed the day 14 survey but not the day 21 survey and interview). As preregistered, analyses were conducted on the participants who completed the full study (85% of participants, $N=155$).
Of participants who completed the full study, 86% ($N=133$) also completed the follow-up survey that was sent four weeks after the final survey and interview (completion by condition: chatbot = 65/78; control = 65/77).
@@
Figure 3: Mediation model for desire to socially connect, magnitude of social impact, and anthropomorphism. Paths between variables designated with letters $a$, $b$, and $c$. Path estimates and significance are indicated (* = $p<0.01$, ** = $p<0.001$, ns = non-significant).
@@
We used non-parametric Wilcoxon rank-sum tests to evaluate differences by condition over time for desire to socially connect, loneliness, social health, magnitude of social impact, and directional social impact scores. Significant differences were found between the chatbot and control groups in their desire to socially connect. Participants in the chatbot condition, versus the control condition, reported higher desire to socially connect when tested on day 7 (chatbot $M=1.95$, control $M=1.46$, $p<0.01$) and 14 (chatbot $M=1.94$, control $M=1.49$, $p<0.05$). No difference was found on day 1 or day 21. No difference was found on day 1 or day 21. No significant results emerged for magnitude of social impact, directional social impact, loneliness, or social health. Compared to those who played word games, those who engaged with the companion chatbot did not have significant impacts to their social health, loneliness, or social interactions or relationships, but did show a significant effect on their desire to socially connect with others.
@@
To test the hypothesis that people with a higher desire to socially connect on day 1 would anthropomorphize the chatbot more (H2), we ran non-parametric Spearman’s correlations with Holm correction for multiple comparisons between participants’ desire to socially connect on day 1 and their anthropomorphism of the chatbot over time. Figure 1A shows the day-1 score for the desire to socially connect versus the anthropomorphism scores averaged over time (an average of anthropomorphism scores at day 7, 14, and 21). Desire to socially connect significantly predicted anthropomorphism ($\rho=0.40$, $p<0.01$). Participants with a higher desire to socially connect on day 1 were more likely to anthropomorphize the chatbot across 21 days of use. Figure 1B shows the day-1 scores for the desire to socially connect versus the anthropomorphism scores separated by time point (days 7, 14, and 21). The desire to socially connect on day 1 significantly predicted anthropomorphism at all three time points for the chatbot group (day 7: $\rho=0.34$, $p<0.01$, day 14: $\rho=0.45$, $p<0.001$, day 21: $\rho=0.38$, $p<0.01$). H2 was supported: people with a higher desire to socially connect on day 1 were more likely to anthropomorphize the chatbot on day 7, 14, and 21.
In comparison, for the control group, we did not find a significant relationship between the desire to socially connect on day 1 and anthropomorphism of the word games averaged over days 7, 14, and 21 ($\rho=0.07$, $p=0.57$). Likewise, we found no significant relationship between the desire to socially connect on day 1 and anthropomorphism of the word games when tested separately for days 7, 14, and 21 (day 7: $\rho=0.07$, $p=1.0$, day 14: $\rho=0.04$, $p=1.0$, day 21: $\rho=0.07$, $p=1.0$).
@@
In H3, people who anthropomorphized the chatbot more should report that the interaction with the chatbot had a greater social impact on them. To test this hypothesis, we ran non-parametric Spearman’s correlations with Holm correction for multiple comparisons between the participants’ anthropomorphism of the chatbot and the magnitude of social impacts over time. The magnitude of social impact score captures the strength of perceived impact regardless of whether the impact was positive or negative (see Methods).
@@
Figure 2A shows the anthropomorphism scores (averaged over days 7, 14, and 21) versus the social impact scores (averaged over days 7, 14, and 21) for the chatbot group. Anthropomorphism significantly predicted social impact ($\rho=0.52$, $p<0.0001$). People who anthropomorphized the chatbot more indicated that the human-chatbot interaction had a greater impact on their social interactions and relationships with family and friends. Figure 2B shows three regression lines for the chatbot group. The first shows the anthropomorphism score as measured on day 7 versus the social impact score as measured on day 7; the second shows anthropomorphism versus social impact, both measured on day 14; and the third shows anthropomorphism versus social impact, both measured on day 21. For all three regression lines, anthropomorphism significantly predicted social impact (day 7: $\rho=0.43$, $p<0.0001$, day 14: $\rho=0.48$, $p<0.0001$, day 21: $\rho=0.52$, $p<0.0001$). H3 was supported: higher anthropomorphism of the chatbot correlated with a higher magnitude of social impact of the chatbot.
For comparison, in the control group, the relationship between anthropomorphism (averaged over days 7, 14, and 21) and social impact of the word games (averaged over days 7, 14, and 21) was also significant ($\rho=0.33$, $p<0.01$).Those who tended to anthropomorphize the word games also tended to report that the games had a greater social impact on them. When we tested days 7, 14, and 21 separately, we found that the only significant relationship between anthropomorphism and social impact, for the control group, occurred on day 7 ($\rho=0.35$, $p<0.01$). The relationship was not significant for day 14 ($\rho=0.20$, $p=0.09$) or day 21 ($\rho=0.23$, $p=0.09$).
@@
The foregoing analysis considered only the magnitude of social impacts, not the direction. Did higher anthropomorphism of the chatbot correlate specifically with more positive perceived social impacts, more negative perceived impacts, or did different people tend to swing different directions? We performed an additional analysis using the directional social impact measure. (The individual, raw scores for the two social impact questions were averaged, resulting in a scale that ranged from positive to negative values; see Methods.) We found that higher anthropomorphism predicted more positive social impacts at all time points for the chatbot group (day 7: $\rho=0.61$, $p<0.0001$, day 14: $\rho=0.67$, $p<0.0001$, day 21: $\rho=0.71$, $p<0.0001$).People who anthropomorphized the chatbot more were more likely to report positive impacts to their social interactions and relationships with family and friends. For comparison, for the control group, a significant positive relationship between anthropomorphism and social impact occurred on day 7 ($\rho=0.35$, $p<0.01$) but was not significant for day 14 ($\rho=0.20$, $p=0.18$) or day 21 ($\rho=0.16$, $p=0.18$).
@@
In addition to asking how participants perceived the impact of the chatbot on their human relationships, we also tested social health through a separate scale. According to H4, people who anthropomorphized the chatbot more should show greater changes to their social health scores. To test this hypothesis, we ran Spearman’s correlations with Holm correction for multiple comparisons. We tested the relationship between the mean anthropomorphism scores for each participant (averaged over days 7, 14, and 21) and the mean social health scores (averaged over days 7, 14, and 21). We also tested the relationship between the anthropomorphism score and the social health score for each of the three time points individually. We found no significant relationships for the chatbot group ($p>0.05$). H4 was not supported. We also did not find significant relationships for the corresponding analyses in the control group..
@@
According to H5, the degree of anthropomorphism of the chatbot should mediate the relationship between people’s desire to socially connect and the magnitude of social impact of interacting with the chatbot. For this analysis, as preregistered, we did not evaluate social health impacts, as those relationships were non-significant in prior analyses. Instead, we evaluated the scores on perceived social impact. We ran a mediation model with bootstrapped standard errors (1000 resamples) using the lavaan package (Rosseel 2012). Figure 3 depicts the output of the mediation model, including the paths estimates and significance values.
For the chatbot group, results supported a mediation relationship. We found a significant positive association between the desire to socially connect and anthropomorphism ($a=0.54$, $p<0.001$) and between anthropomorphism and social impacts ($b=0.30$, $p<0.01$). The indirect effect of desire to socially connect, via anthropomorphism, on social impacts, was significant ($ab=0.16$, $p<0.05$, 95% CI: [0.03, 0.33]). In contrast, the direct effect of desire to socially connect on social impacts was not significant ($c=0.13$, $p=0.27$). The total effect was significant ($total=0.29$, $p<0.01$, 95% CI: [0.05, 0.48]), providing evidence for partial mediation. Approximately 57% of the total effect was explained by the indirect effect of anthropomorphism. The remaining 43% of the total effect was not explained by anthropomorphism and may be due to direct effects or other unmeasured mediators. These findings suggest that anthropomorphism plays a significant role in linking people’s desire to socially connect to their reported social impacts of interacting with the chatbot. H5 was supported.
In comparison, the results for the control condition indicated that the association between the desire to socially connect and anthropomorphism was not significant ($a=0.08$, $p=0.58$). The association between anthropomorphism and magnitude of social impacts was significant ($b=0.31$, $p<0.01$). The indirect effect was not significant ($ab=0.02$, $p=0.63$), nor was the direct effect of desire to socially connect on magnitude of social impacts ($c=0.16$, $p=0.05$). The total effect was also non-significant ($total=0.19$, $p=0.09$). These findings suggest that the mediating role of anthropomorphism is specific to the chatbot condition.
@@
The stepwise linear regression model (AIC) included 20 predictors and explained 75% of the variance in anthropomorphism of the chatbot (adj $R^{2}=0.75$, $AIC=146.1$, $p<0.0001$). Significant predictors within the model with positive estimates were the desire to socially connect ($b=0.42$, $p<0.001$), beliefs about AI agency ($b=0.21$, $p<0.01$), the General Attitudes Toward Robots Scale (GAToRS) personal-positive scale (P+) ($b=0.23$, $p<0.05$), social health ($b=1.10$, $p<0.001$), age ($b=0.04$, $p<0.01$), mental health vulnerability ($b=0.34$, $p<0.001$), and identifying as having multiple sexual orientations ($b=0.77$, $p<0.05$). Significant predictors within the model with negative estimates were beliefs about AI experience (b=0.36b=-0.36, $p<0.01$), GAToRS society-negative scale (S-) (b=0.25b=-0.25, $p<0.05$), extraversion (b=0.22b=-0.22, $p<0.01$), self-esteem (b=0.38b=-0.38, $p<0.01$), social competence (b=0.33b=-0.33, $p<0.05$), and identifying as lesbian (b=1.13b=-1.13, $p<0.05$) or the sexual orientation of “prefer to self-describe” (b=2.27b=-2.27, $p<0.01$).
@@
We then compared the magnitude of social impact scores across time and at each time point between the vulnerable and non-vulnerable subgroups using Wilcoxon rank-sum tests. We found no significant differences across time for the magnitude of social impact (averaged scores for day 7, 14, and 21) between vulnerable and non-vulnerable subgroups. More vulnerable participants did not report greater social impacts than the non-vulnerable group ($p=0.40$). We then evaluated differences at each time point using Holm’s correction for multiple comparisons and did not find significant differences in magnitude of social impact scores between subgroups. Participants in the vulnerable group did not report greater social impacts than those in the non-vulnerable group on day 7, 14, or 21 ($p$ adj for each comparison >0.05>0.05). H6 was not supported. For comparison, in the control group, participants in the vulnerable group also did not report greater social impacts than those in the non-vulnerable group ($p$ adj for each comparison >0.05>0.05).
Attachment to Companion Chatbots
We tested the prediction that the chatbot would not be more habit-forming than the word games (H7). We operationalized habit formation by two metrics measured in the follow-up survey, four weeks after the end of the study. First, we used participants’ report of whether or not they continued to engage in their assigned daily task after the last day of the study (day 21). We used a binary variable (yes or no). Second, we used the reported duration of time spent engaging in the daily task after the last day of the study. Participants’ responses on duration were binned into four categories: a few days, 1-2 weeks, 3-4 weeks, or ongoing.
The number of participants who completed the follow-up survey was not statistically different between the chatbot and control conditions ($\chi^{2}=0$, $p=1$; chatbot: 65 completed, 13 did not; control: 65 completed, 12 did not).
@@
We then evaluated the duration of post-study engagement in the daily task by condition. Of the 14 people who indicated they engaged with the chatbot after the study ended, 6 (43%) indicated that their engagement was ongoing, none (0%) said they used it for 3-4 weeks, 3 (21%) said they used it for 1-2 weeks, and 5 (36%) said they used it for a few days. Of the 40 people who indicated they engaged with the word games after the study ended, 26 (65%) indicated that their engagement was ongoing, 3 (8%) said they used it for 3-4 weeks, 6 (15%) said they used it for 1-2 weeks, and 5 (13%) said they used it for a few days. A higher percentage of people in the control condition reported the longest post-task duration of engagement (ongoing engagement) (65%) as compared to the chatbot condition (43%). However, these ratios were not significantly different from one another from a chi-squared approximation with continuity correction ($\chi^{2}=1.3$, $p=0.26$).
H7 was partially supported. The chatbot was not more habit-forming than the word games. On the contrary, people who played the word games were more likely to engage in that task after the study had ended, as compared to people who interacted with the chatbot.